Winston Clinic is evaluating a project that costs $52,125 and has expected net cash inflows of $12,000 per year for eight years. The first inflow occurs one year after the cost outflow, and the project has a cost of capital of 12 percent.


 a. What is the project's payback?

 b. What is the project's NPV? Its IRR? Its MIRR?

 c. Is the project financially acceptable? Explain your answer
Answer #1

a)

Payback of project = 52,125 / 12,000 = 4.34 years

b)

NPV = present value of cash inflows - present value of cash outflows

Present value of cash inflows = annuity * [ 1 - 1 / ( 1 + r)n]] / r

Present value of cash inflows = 12,000 * [ 1 - 1 / ( 1 + 0.12)8]] / 0.12

Present value of cash inflows = 12,000 * 4.96764

Present value of cash inflows = 59,611.6772

NPV = 59,611.6772 - 52,125

NPV = $7,486.68

IRR is the rate of return that makes NPV equal to 0

52,125 - 12,000 * [ 1 - 1 / ( 1 + R)8]] / R = 0

Using trial and error method, i.e, after trying various values for R, lets try 16

52,125 - 12,000 * [ 1 - 1 / ( 1 + 0.16)8]] / 0.16 = 0

0 = 0

Therefore, IRR is 16%

Future value of 12,000 = annuity * [( 1 + r)n - 1] / r

Future value of 12,000 = 12,000 * [( 1 + 0.12)8 - 1] / 0.12

Future value of 12,000 = 12,000 * 12.299693

Future value of 12,000 = 147,596.317

Future value = present value ( 1 + r)n

147,596.317 = 52,125 ( 1 + r)8

2.83158 = ( 1 + r)8

8th root of 2.83158 is 1.138947

1.138947 = 1 + r

R = 0.138947 or 13.895%

MIRR is 13.895%

The project if financially acceptable as it has a positive NPV and IRR & MIRR are greater than cost of capital of 12%.